OpenStudy (anonymous):

Solve the following equation: 6n^2=5n

OpenStudy (jim766):

first make it equal to zero subtract 5n on both sides

OpenStudy (anonymous):

subtract 5n?

OpenStudy (jim766):

you end up with 6n^2 - 5n = 0 now factor it, what is GCF?

OpenStudy (anonymous):

omg!! no no no I meant 6n^2=15n my bad

OpenStudy (jim766):

6n^2 = 5n -5n -5n 6n^2 - 5n = 0 now factor it n(6n - 5) = 0 now use zero product property

OpenStudy (anonymous):

where do you keep getting 5???

OpenStudy (anonymous):

its suppose to 15

OpenStudy (jim766):

sorry 6n^2 -15n = 0 GCf 3n(2n - 5) = 0 now use zero prod property

OpenStudy (anonymous):

x=0,5/2